The demand function for a Christmas music CD is given by

q=D(p)=0.25(225−p2)

where qq (measured in units of a hundred) is the quantity demanded per week and pp is the unit price in dollars.

(a) Find the elasticity function E(p)=  

(b) Evaluate the elasticity at 10. E(10)=  

(c) Should the unit price be lowered slightly from 10 in order to increase revenue?   ?    yes    no   

(d) Use the elasticity of demand to find the price which maximizes revenue for this product. p= dollars

Round your answer to two decimal places as needed.
Answer #1

Griuen a= 0.25(225-p bierenhiati eauation C uor.t get 0-0.25x2 P day -0.SP Formulas for elasficity is ELPI- da dp SupshituteDate Jomala for elatkeity gel, ELP) -O.5 x P =0Spx P (56-25-0.25p2) ECP) -(-0.5) x p2 (56-25-0-25p2) (10) (56-25-0-25x10o) (ECIO)-16 maams tuat a 1 decreaue in price belous 10 iu imureaue theauanhty domanded 6% Thare will be tuo elacts Fixatly SezaNo. Date d(TR) dp う 4 + P.das dp ニ0 Pda 1 + ECP)0 tu valua o E (P) from Subshitute part we I (-0.5)x P 56-25-0-25p) 2 1 = (56Date 56.25 75 +53 うP: 1 Sinu bria cannot be migabu, tas means Pz 53 s the greuenu triu maximising 8.66